ImSonique:

What is the value of y in the equation 5x + 2y = 20, when x = 0.3? 2.5 2.8 9.25 10.75

Elephant2020:

What do you think it is?

Elephant2020:

Yes

Elephant2020:

That's the answer

ImSonique:

thank u so much

Xaophu:

Step 1: Simplify both sides of the equation. (5)(0.3)+2y=20 1.5+2y=20 2y+1.5=20 Step 2: Subtract 1.5 from both sides. 2y+1.5−1.5=20−1.5 2y=18.5 so y=9.25

Xaophu:

wait step 3 is Divide both sides by 2.

Elephant2020:

Yes you got it

Xaophu:

here if you ever get confused try using this, it gives you step by step answers too https://www.mathpapa.com/algebra-calculator.html

Xaophu:

it's a really helpful tool for algebraic equations

Elephant2020:

Yes